OpenStudy (muddygirl):

Use the distributive property to rewrite the expression and solve. 8*63 = 8*__ + 8*3=63

OpenStudy (plasmataco):

The distrubutive property is a(b+c)=ab+ac

OpenStudy (welshfella):

how did you get 21?

OpenStudy (plasmataco):

Plug 63 into (b+c)

OpenStudy (plasmataco):

8 into a

OpenStudy (plasmataco):

now you know that c is 3

OpenStudy (muddygirl):

The distributive property is a(b+c)=a*b+a*c and a(b-c)=a*b-a*c

OpenStudy (plasmataco):

so you get 8*b+8*3=8*63

OpenStudy (plasmataco):

8*63 is 504

OpenStudy (welshfella):

- so b = 60

OpenStudy (plasmataco):

8*3 is 24

OpenStudy (plasmataco):

504-24=480

OpenStudy (plasmataco):

480/8=6 yes, sry I said 24 in the beginning

OpenStudy (muddygirl):

Ok but i have to rewrite and solve for the blank space

OpenStudy (plasmataco):

so plug 60 into the blank space taylor swift style (sry I had to)

OpenStudy (plasmataco):

XD

OpenStudy (muddygirl):

i did to lol

OpenStudy (plasmataco):

All you gotta do

OpenStudy (plasmataco):

is have some fun

OpenStudy (plasmataco):

XD I'm so so sry

OpenStudy (welshfella):

ONe question - where does the '=63' come from ? thats not correct

OpenStudy (plasmataco):

The beginning, ignore the last =63

OpenStudy (welshfella):

yea I did. But it is confusing

OpenStudy (triciaal):